cosecx-sinx=m and secx-cosx=n eliminate x

Grade:12th pass

1 Answers

Sher Mohammad IIT Delhi
10 years ago
1-(sinx)^2=msinx
1-(cosx)^2=ncosx

add both terms
1-((cosx)^2+(sinx)^2)=msinx+ncosx
1-1=msinx+ncosx
msinx=-ncosx
tanx=-n/m
x=inverse(tan)(-n/m)

sher mohammad
b.tech, iitd